OpenStudy (anonymous):

3x - 17 = 5y 4y + 18 = -2x Help me set up these equations so I can solve them. I just need the first step.

OpenStudy (anonymous):

@zepdrix

zepdrix (zepdrix):

Still having trouble moving the variables around? :o

OpenStudy (anonymous):

Yes :( I worked it all out and got a decimal, which isn't an answer choice.

zepdrix (zepdrix):

\[\Large\rm 3x-17=5y\]Subtracting 5y from each side gives us:\[\Large\rm 3x-17\color{red}{-5y}=5y\color{red}{-5y}\]So that cancels out the 5y's on the right side,\[\Large\rm 3x-17-5y=0\]Adding 17 to each side gives us:\[\Large\rm 3x-17\color{orangered}{+17}-5y=0\color{orangered}{+17}\]17's cancel out on the left,\[\Large\rm 3x-5y=17\]

OpenStudy (anonymous):

Okay I have that.

OpenStudy (anonymous):

Now I'm at 4y + 2x = -18

zepdrix (zepdrix):

Hmm ok let's see what went wrong then.\[\Large\rm 3x-5y=17\]\[\Large\rm \color{royalblue}{4y + 18 = -2x}\]Oh we have to mess with the second equation also right?\[\Large\rm 3x-5y=17\]\[\Large\rm \color{royalblue}{2x+4y= -18}\]

zepdrix (zepdrix):

Yah looks like you've got it setup correctly.

OpenStudy (anonymous):

So should I multiply the top by 2 and the bottom by -2?

OpenStudy (anonymous):

To cancel out the x's of course

zepdrix (zepdrix):

The `Least Common Multiple` of 3 and 2 is 6. (This is assuming that we want to cancel out the x's). So we'd multiply the top one by 2. And the bottom one by -3.

OpenStudy (anonymous):

Yep, yet another typo. That's what I meant. Let me try it now

OpenStudy (anonymous):

I'm at -6x + -12y = 54 and 6x - 10y = 34

zepdrix (zepdrix):

Mmmm seems correct so far.

OpenStudy (anonymous):

I finally got it. Thank you so much for all your help. I'm in a class with no friends so there's no one else I can ask for help.

zepdrix (zepdrix):

No problem c: Mmm that sucks!